OpenStudy (xguardians):

MEDAL AND FAN, PLEASE EXPLAIN ENTIRE WALKTHROUGH! 1. Lead will react with hydrochloric acid to produce lead (ii) chloride and hydrogen. How many moles of hydrochloric acid are needed to completely react with 0.36 mol of lead?

OpenStudy (xguardians):

@SolomonZelman , @mathstudent55 , @aaronq , @radar

OpenStudy (xguardians):

@ganeshie8 , @Nurali Someone please help me?

rishavraj (rishavraj):

0.72 moles

OpenStudy (xguardians):

Can you please explain how you got that, (preferably the setup_ @rishavraj

rishavraj (rishavraj):

[Pb + 2HCl \rightarrow PbCl _{2} + H _{2}\]

rishavraj (rishavraj):

was that enough..

OpenStudy (xguardians):

Thanks, yeah I figured it out on my own. But your answer helped me confirm that I was right!

rishavraj (rishavraj):

anytime....

OpenStudy (aaronq):

I know the question is answered, but i want to clarify you arrive at that. you wanna use a ratio of the moles and stoichiometric coefficients of the reactants: \(\dfrac{n_{HCl}}{2}=\dfrac{n_{Pb}}{1}\rightarrow n_{HCl}=\dfrac{2*n_{Pb}}{1}\) (recall n = moles) This confirms that you need twice as many moles to completely react.